### v1

@@ 4.3 Resampling Trick
The proof is presented in Appendix 8.6. Proposition 5 indicates that the resampling technique serves to “rectify” the distributional divergence caused by guidance. This aspect is crucial, as we have demonstrated that training-free guidance is particularly susceptible to adversarial gradients, which steer the images away from the natural image distribution. The resampling technique realigns these images with the distribution of natural images. This theoretical insight is corroborated by the empirical evidence presented in Figure 2 of [44].
5 Experiments
In this section, we evaluate the efficacy of our proposed techniques across various diffusion models, including CeleA-HQ, ImageNet, and text-to-motion, under different conditions such as sketch, segmentation, text, and object avoidance. We compare our methods with established baselines: Universal Guidance (UG) [2], Loss-Guided Diffusion with Monte Carlo (LGD-MC) [33], and Training-Free Energy-Guided Diffusion Models (FreeDoM) [44]. UG employs guidance as delineated in (5) and uses resampling strategies outlined in Algorithm 3 to enhance performance. Similarly, FreeDoM is founded on (5) and resampling. In addition, FreeDoM incorporates a time-dependent step size for each gradient guidance and judiciously selects the diffusion step for executing resampling. LGD-MC also utilizes guidance from (6), setting the $n=10$. For the sampling process, we employ the DDIM method with 100 steps, consistent with the protocols in FreeDoM and LGD-MC papers. We have developed our approach based on FreeDoM, with the inclusion of random augmentation (Algorithm 1) and projected gradient descent (Algorithm 2).
5.1 Guidance to CelebA-HQ Diffusion
For benchmarking, we preprocess the CelebA-HQ dataset with BiSeNet to extract segmentation maps and use the model checkpoint from [41] to acquire the sketches. We randomly select $1000$ samples each of segmentation maps, sketches, and text descriptions. The comparative results are presented in Table 1. Consistent with [44], the resampling number for both FreeDoM and UG is set to $s=1$. Figure 4 displays a random selection of the generated images. More image samples are provided in the supplementary materials.
Methods Segmentation maps Sketches Texts
Distance↓ FID↓ Distance↓ FID↓ Distance↓ FID↓
UG [2] 2247.2 39.91 52.15 47.20 12.08 44.27
LGD-MC [33] 2088.5 38.99 49.46 54.47 11.84 41.74
FreeDoM [44] 1657.0 38.65 34.21 52.18 11.17 46.13
Ours 1575.7 33.31 30.41 41.26 10.72 41.25
5.2 Guidance to ImageNet Diffusion
Methods LGD-MC [33] UG [2] FreeDoM [44] Ours
CLIP Score↑ 24.3 25.7 25.9 27.7
For the unconditional ImageNet diffusion, we employ text guidance in line with the approach used in FreeDoM and UG [2, 44]. We utilize CLIP-B/16 as the image and text encoder, with cosine similarity serving as the loss function to measure the congruence between the image and text embeddings. To evaluate performance and mitigate the potential for high-scoring adversarial images, we use CLIP-L/14 for computing the CLIP score. In FreeDoM, resampling is conducted for time steps ranging from $800$ to $300$, with the number of resamplings fixed at $10$, as described in [44]. Given that UG resamples at every step, we adjust its resampling number $s=5$ to align the execution time with that of FreeDoM. The textual prompts for our experiments are sourced from [10, 21]. The comparison of different methods is depicted in Table 2. The corresponding randomly selected images are illustrated in Figure 5. The table indicates that our method achieves the highest consistency with the provided prompts. As shown in Figure 5, LGD-MC tends to overlook elements of the prompts. Both UG and FreeDoM occasionally produce poorly shaped animals, likely influenced by adversarial gradients. Our approach addresses this issue through the implementation of random augmentation. Additionally, none of the methods successfully generate images that accurately adhere to positional prompts such as “left to” or “below”. This limitation is inherent to CLIP and extends to all text-to-image generative models [37]. More image samples are provided in the supplementary materials.
5.3 Guidance to Human Motion Diffusion
Methods “Backwards” “Balanced Beam” “Walking” “Jogging”
Loss↓ CLIP↑ Loss↓ CLIP↑ Loss↓ CLIP↑ Loss↓ CLIP↑
In this subsection, we extend our evaluation to human motion generation using the Motion Diffusion Model (MDM) [36], which represents motion through a sequence of joint coordinates and is trained on a large corpus of text-motion pairs with classifier-free guidance. We apply the object avoidance guidance as described in [33]. Let $\bm{x}_{0}(t)$ denote the joint coordinates at time $t$, $\bm{y}_{t}$ the target location, $\bm{y}_{\text{obs}}$ the obstacle location, $r$ the radius of the objects, and $T$ the total number of frames. The loss function is defined as follows:
Our experimental configuration adheres to the guidelines set forth in [33]. We assess the methods using the targeting loss (the first term in (9)), the object avoidance loss (the second term in (9)), and the CLIP score calculated by MotionCLIP [35]. For FreeDoM, the number of resamplings is set to $3$. In our method, random augmentation is omitted because the guidance is not computed by neural networks so the adversarial gradient issues are absent. The quantitative results of our investigation are summarized in Table 3, while Figure 6 showcases randomly selected samples. Our methods exhibit enhanced control quality over the generated motion. The videos are provided in the supplementary materials.
6 Conclusions
In this paper, we conducted a comprehensive investigation into training-free guidance, which employs pre-trained diffusion models and guides them using the off-the-shelf trained on clean images. Our exploration delved into the underlying mechanisms and fundamental limits of these models. Moreover, we proposed a set of enhancement techniques and verified their effectiveness both theoretically and empirically.
Limitations. Despite our efforts to mitigate the shortcomings of training-free methods and enhance their performance, certain intractable limitations remain. Notably, the refined training-free guidance still necessitates a higher number of NFEs when compared with extensive training methods such as classifier-free guidance. This is because adversarial gradient cannot be fully eliminated without training.
Ethical Consideration. Similar to other models designed for image creation, our model also has the unfortunate potential to be used for creating deceitful or damaging material. We pledge to restrict the usage of our model exclusively to the realm of research to prevent such misuse.

### v2

@@ C.2 Time-travel Trick
The proof is presented in Appendix D.6. Proposition C.3 indicates that the time-travel technique serves to “rectify” the distributional divergence caused by guidance. This aspect is crucial, as we have demonstrated that training-free guidance is particularly susceptible to adversarial gradients, which steer the images away from the natural image distribution. The time-travel technique realigns these images with the distribution of natural images.
C.3 Efficiency of Random Augmentation
Given the multiple invocations of the guidance network necessitated by random augmentation (RA), concerns regarding the efficiency of this approach are understandable. However, it is important to note that, compared to the diffusion backbone, the guidance network exhibits a more lightweight architecture, thereby mitigating any significant increase in computational demand. To empirically illustrate this point, we present the computation times associated with varying degrees of augmentation in Table 4. In the conducted experiments, we set the cardinality of the set $\mathcal{T}$ to $10$, thereby having little impact on the inference time. These experiments were conducted on a single NVIDIA A100 GPU.
Setting w/o RA $|\mathcal{T}|=1$ $|\mathcal{T}|=10$ $|\mathcal{T}|=20$ $|\mathcal{T}|=30$
CLIP Score 0.541 0.544 0.571 0.592 0.625
Table 4: Inference time (seconds per diffusion step) of different random augmentation configurations. The diffusion backbone is ImageNet diffusion and the guidance network is CLIP-B/16.
Appendix D Proofs
D.1 Definitions
This subsection introduces a few definitions that are useful in the following sections.
Definition D.1.
($L$-Lipschitz) A function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is said to be $L$-Lipschitz if there exists a constant $L\geq 0$ such that $\|f(\bm{x}_{2})-f(\bm{x}_{2})\|\leq L\|\bm{x}_{2}-\bm{x}_{1}\|$ for all $\bm{x}_{1},\bm{x}_{2}\in\mathbb{R}^{n}$.
Definition D.2.
(PL condition) A function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ satisfies PL condition with parameter $\mu$ if $\|\nabla f(\bm{x})\|^{2}\geq\mu f(\bm{x})$.
Definition D.3.
(Q function) We define the Gaussian tail probability $Q(x)=\mathbb{P}(x\geq a)$, $x\sim\mathcal{N}(0,1)$.
D.2 Proof for Proposition 3.1
Denote $\hat{\bm{x}}_{0}=\mathbb{E}_{p(\bm{x}_{0}|\bm{x}_{t})}(\bm{x}_{0})$, the gradient guidance term in (5) can be written as the following:
where the last equality follows the second-order Miyasawa relationship $\text{Cov}[\bm{x}_{0}|\bm{x}_{t}]=\sigma_{t}^{2}(\bm{I}+\sigma_{t}^{2}\nabla^{2}\log p_{t}(\bm{x}_{t}))$ [28].
For the first condition, the Lipschitz constant satisfies
$|\ell_{t}(\bm{x}_{1})-\ell_{t}(\bm{x}_{2})|$ $\leq\frac{L_{f}}{\sqrt{\alpha_{t}}}|\bm{x}_{1}-\nabla_{\bm{x}_{1}}\log p_{t}(\bm{x}_{1})-\bm{x}_{2}+\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{2})|$
$\leq L_{f}(1+L_{p})\|\bm{x}_{1}-\bm{x}_{2}\|,$
and the PL constant satisfies
$\ell_{t}(\bm{x}_{t})\leq\frac{1}{\mu}\left\|\frac{\partial\ell}{\partial\hat{\bm{x}}_{0}}\right\|^{2}=\frac{1}{\mu}\|\nabla_{\bm{x}_{t}}\ell\cdot\text{Cov}^{-1}(\bm{x}_{0}|\bm{x}_{t})\sigma_{t}^{2}\sqrt{\alpha_{t}}\|^{2}\leq\frac{\sigma_{t}^{4}\alpha_{t}}{\mu\lambda_{min}^{2}}\|\nabla_{\bm{x}_{t}}\ell\|^{2}.$
The second and third conditions directly follow Lemma D.4.
Lemma D.4.
(Linear Convergence Under PL condition; [20]) Denote $\bm{x}^{0}$ as the initial point and $\bm{x}^{t}$ as the point after $t$ gradient steps. If the function is $L$-Lipschitz and $\mu$-PL, gradient descent with a step size $\eta=\frac{1}{L}$ converges to a global solution with $\ell(\bm{x}^{t})\leq(1-\mu/L)^{t}\ell(\bm{x}^{0})$.
D.3 Proof for Proposition 3.2
Proof.
The proof of this theorem is based on the proof of Lemma 1 in [32]. By the definition of expectation, we have
$\hat{f}(\bm{x})$ $=\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})}[f(\bm{x}+\sigma_{t}\bm{\epsilon})]=(f\circledast\mathcal{N}(0,\sigma_{t}^{2}\bm{I}))(\bm{x})$
$=\frac{1}{(2\pi\sigma_{t}^{2})^{n/2}}\int_{\mathbb{R}^{n}}f(\bm{z})\exp\left(-\frac{1}{2\sigma_{t}^{2}}\|\bm{x}-\bm{z}\|^{2}\right){\rm d}\bm{z}.$
We then show that for any unit direction $\bm{u}$, $\bm{u}^{T}\nabla\hat{f}(\bm{x})\leq\sqrt{\frac{2}{\pi\sigma_{t}^{2}}}$. The derivative of $\hat{f}$ is given by
$\nabla\hat{f}(\bm{x})$ $=\frac{1}{(2\pi\sigma_{t}^{2})^{n/2}}\int_{\mathbb{R}^{n}}f(\bm{z})\nabla\exp\left(-\frac{1}{2\sigma_{t}^{2}}\|\bm{x}-\bm{z}\|^{2}\right){\rm d}\bm{z}$
$=\frac{1}{(2\pi\sigma_{t}^{2})^{n/2}\sigma_{t}^{2}}\int_{\mathbb{R}^{n}}f(\bm{z})(\bm{x}-\bm{z})\exp\left(-\frac{1}{2\sigma_{t}^{2}}\|\bm{x}-\bm{z}\|^{2}\right){\rm d}\bm{z}.$
Thus, the Lipschitz constant is computed as
$\bm{u}^{T}\nabla\hat{f}(\bm{x})$ $\leq\frac{C}{(2\pi\sigma_{t}^{2})^{n/2}}\int_{\mathbb{R}^{n}}|\bm{u}^{T}(\bm{x}-\bm{z})/\sigma_{t}^{2}|\exp\left(-\frac{1}{2\sigma_{t}^{2}}\|\bm{x}-\bm{z}\|^{2}\right){\rm d}\bm{z}$
$=\frac{C}{(2\pi\sigma_{t}^{2})^{1/2}}\int_{-\infty}^{\infty}|s|\exp\left(-\frac{1}{2}s^{2}\right){\rm d}s=\sqrt{\frac{2}{\pi\sigma_{t}^{2}}}.$
Similarly, for the Lipschitz constant of the gradient, we have
$\|\nabla^{2}\hat{f}(\bm{x})\|_{\text{op}}\leq\|\nabla^{2}\hat{f}(\bm{x})\|_{\text{2}}$
$\leq\frac{C}{\sqrt{2\pi}\sigma_{t}\cdot\sigma_{t}^{4}}\left(\int_{-\infty}^{\infty}s^{2}\exp(-\frac{1}{2}s^{2}/\sigma_{t}^{2}){\rm d}s+\int_{-\infty}^{\infty}\sigma_{t}^{2}\exp(-\frac{1}{2}s^{2}/\sigma_{t}^{2}){\rm d}s\right)=\frac{2C}{\sigma_{t}}.$
D.4 Proof for Proposition 3.3
Proof.
We first analyze the discretization error of a single step from time $s$ to $t$. We denote the update variable for DDIM is $\bm{x}_{t}^{*}$ and the optimal solution of the diffusion ODE at time $t$ as $\bm{x}_{t}^{*}$. Let $\sigma_{t}=\sqrt{1-\alpha_{t}}$, $\lambda_{t}=\frac{1}{2}\log(\frac{\alpha_{t}}{1-\alpha_{t}})$ and $h=\lambda_{t}-\lambda_{s}$. According to (B.4) of [24], the update of DDIM solver is given by
$\bm{x}_{t}=\sqrt{\frac{\alpha_{t}}{\alpha_{s}}}\bm{x}_{s}-\sigma_{t}(e^{h}-1)\bm{\epsilon}_{\theta}(\bm{x}_{s},s).$
A similar relationship can be obtained for the optimal solution
$\bm{x}_{t}^{*}=\sqrt{\frac{\alpha_{t}}{\alpha_{s}}}\bm{x}_{s}^{*}-\sigma_{t}(e^{h}-1)\bm{\epsilon}_{\theta}(\bm{x}_{s}^{*},s)+O(h^{2}).$
We then bound the error between $\bm{x}_{t}$ and $\bm{x}_{t}^{*}$. We have
$\bm{x}_{t}$ $\leq\sqrt{\frac{\alpha_{t}}{\alpha_{s}}}\bm{x}_{s}-\sigma_{t}(e^{h}-1)(\bm{\epsilon}_{\theta}(\bm{x}_{s}^{*},s)+L(\bm{x}_{s}-\bm{x}_{s}^{*}))$
$=\sqrt{\frac{\alpha_{t}}{\alpha_{s}}}\bm{x}_{s}^{*}-\sigma_{t}(e^{h}-1)\bm{\epsilon}_{\theta}(\bm{x}_{s}^{*},s)+\sqrt{\frac{\alpha_{t}}{\alpha_{s}}}(\bm{x}_{t}-\bm{x}_{s}^{*})-\sigma_{t}(e^{h}-1)L(\bm{x}_{s}-\bm{x}_{s}^{*}))$
$=\bm{x}_{t}^{*}+O(h^{2})+O(Lh^{2}).$
If we run DDIM for $M$ steps, $h_{\max}=O(1/M)$, and we achieve the discretization error bound for DDIM algorithm:
$\bm{x}_{0}=\bm{x}^{*}_{0}+O(M(h_{\max}^{2}+h_{\max}^{2}L))=\bm{x}^{*}_{0}+O(h_{\max}+h_{\max}L).$
D.5 Proof for Proposition 4.1
Proof.
By the definition of expectation, we have
$\hat{f}(\bm{x})$ $=\mathbb{E}_{\bm{\epsilon}\sim p(\bm{\epsilon})}[f(\bm{x}+\sigma_{t}\bm{\epsilon})]=(f\circledast p)(\bm{x})=\int_{\mathbb{R}^{n}}f(\bm{z})p(\bm{x}-\bm{z}){\rm d}z.$
Then we compute the Lipschitz constant
$\bm{u}^{T}\nabla\hat{f}(\bm{x})\leq C\int_{\mathbb{R}^{n}}\|\nabla p(\bm{x}-\bm{z})\|_{2}{\rm d}\bm{z}=C\int_{\mathbb{R}^{n}}\|\nabla p(\bm{z})\|_{2}{\rm d}\bm{z}.$
As for the gradient Lipschitz constant, we have
$L=\|\nabla^{2}\hat{f}(\bm{x})\|_{\text{op}}\leq C\left\|\int_{\mathbb{R}^{n}}\nabla^{2}p(\bm{x}-\bm{z}){\rm d}\bm{z}\right\|_{\text{op}}\leq C\int_{\mathbb{R}^{n}}\|\nabla^{2}p(\bm{z})\|_{\text{op}}{\rm d}\bm{z},$
where $\|\cdot\|_{\text{op}}$ is the operator norm of a matrix. ∎
D.6 Proof for Proposition C.3
Our proof is based on the proof for Theorem 2 in [44]. The total error can be decomposed as the discrepancy between $q(\bm{x}^{1}_{t})$ and $p_{t}(\bm{x}_{t})$ and the error introduced by the ODE solver. We introduce an auxiliary process $q^{r}_{t}(\bm{y}_{t}^{i})$, which initiates from the true distribution and perform the time-travel steps. By triangle inequality:
$\text{TV}(q_{t}(\bm{x}_{t}^{i}),p_{t}(\bm{x}_{t}))\leq\text{TV}(q_{t}(\bm{x}_{t}^{i}),q^{r}_{t}(\bm{y}_{t}^{i}))+\text{TV}(q^{r}_{t}(\bm{y}_{t}^{i}),p_{t}(\bm{x}_{t})),$
where the second term of RHS is bounded by $O(h_{\max})$ according to Proposition 3.3. Then we follow Lemma 1 in [44] to bound the first term.
$\mathbb{E}[\bm{1}[\bm{x}_{t}^{i+1}\neq\bm{y}_{t}^{i+1}]]\leq\mathbb{E}[\bm{1}[\bm{x}_{t+1}^{i+1}\neq\bm{y}_{t+1}^{i+1}]]$
$\overset{(a)}{\leq}$ $\mathbb{E}\left[\left(1-Q\left(\frac{\sqrt{\beta}_{t}B}{2\sqrt{1-\beta_{t}}}\right)\right)\bm{1}[\bm{x}_{t}^{i}\neq\bm{y}_{t}^{i}]\right]$
$=$ $\left(1-Q\left(\frac{\sqrt{\beta}_{t}B}{2\sqrt{1-\beta_{t}}}\right)\right)\mathbb{E}\left[\bm{1}[\bm{x}_{t}^{i}\neq\bm{y}_{t}^{i}]\right],$
where (a) uses Lemma C.2 by setting $\bm{x}_{t+1}^{i+1}=\sqrt{\beta_{t}}\bm{x}_{t}^{i}+\sqrt{1-\beta_{t}}\bm{n}$ and $\bm{y}_{t+1}^{i+1}=\sqrt{\beta_{t}}\bm{y}_{t}^{i}+\sqrt{1-\beta_{t}}\bm{n}$. Applying the bound recursively, we have
$\mathbb{E}[\bm{1}[\bm{x}_{t}^{s}\neq\bm{y}_{t}^{s}]]\leq\left(1-Q\left(\frac{\sqrt{\beta}_{t}B}{2\sqrt{1-\beta_{t}}}\right)\right)^{s-1}\mathbb{E}\left[\bm{1}[\bm{x}_{t}^{1}\neq\bm{y}_{t}^{1}]\right].$
The conclusion follows by $\text{TV}(q_{t}(\bm{x}_{t}^{i}),q^{r}_{t}(\bm{y}_{t}^{i}))\leq\mathbb{P}(\bm{x}_{t}^{s}\neq\bm{y}_{t}^{s})=\mathbb{E}(\bm{1}(\bm{x}_{t}^{s}\neq\bm{y}_{t}^{s}))$ and the initial coupling $\text{TV}(q_{t}(\bm{x}_{t}^{1}),q^{r}_{t}(\bm{y}_{t}^{1}))=\mathbb{P}(\bm{x}_{t}^{1}\neq\bm{y}_{t}^{1})$. Therefore, we conclude that
$\text{TV}(q_{t}(\bm{x}_{t}^{i}),p_{t}(\bm{x}_{t}))\leq\text{TV}(q_{t}(\bm{x}_{t}^{i}),q^{r}_{t}(\bm{y}_{t}^{i}))+\text{TV}(q^{r}_{t}(\bm{y}_{t}^{i}),p_{t}(\bm{x}_{t}))$
$\leq\left(1-Q\left(\frac{\sqrt{\beta}_{t}B}{2\sqrt{1-\beta_{t}}}\right)\right)^{s-1}\text{TV}(q_{t}(\bm{x}_{t}^{1}),q^{r}_{t}(\bm{y}_{t}^{1}))+O(h_{\max})$
$\leq\left(1-Q\left(\frac{\sqrt{\beta}_{t}B}{2\sqrt{1-\beta_{t}}}\right)\right)^{s-1}\text{TV}(q_{t}(\bm{x}_{t}^{1}),p_{t}(\bm{x}_{t}))+O(h_{\max})+O(h_{\max}).$
Appendix E More Qualitative Results
E.1 CelebA-HQ
Figure 7: More qualitative results of CelebA-HQ with zero-shot segmentation guidance. The images are randomly selected.
Figure 8: More qualitative results of CelebA-HQ with zero-shot sketch guidance. The images are randomly selected.
Figure 9: More qualitative results of CelebA-HQ with zero-shot text guidance. The images are randomly selected.
E.2 ImageNet
Figure 10: More qualitative results of ImageNet with zero-shot text guidance. The images are randomly selected.
E.3 Human Motion
See “TeX Source/videos.pptx”.